OpenStudy (anonymous):

a probala has a vertex at (-1,0) and opens down what is the equation

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} y=(x-{\color{brown}{ h}})^2+{\color{blue}{ k}}\\ x=(y-{\color{blue}{ k}})^2+{\color{brown}{ h}} \end{array} \qquad \begin{array}{llll} vertex\ ({\color{brown}{ h}},{\color{blue}{ k}})\\ \end{array}\) bear in mind that, when a parabola opens downwards, it has a leading term with a negative coefficient